Question 4

At December 31, 2016, Bonita Ltd. reported the following as plant assets.
Land £ 2,961,000
Buildings £28,521,000
Less: Accumulated depreciation-buildings 11,791,000 16,730,000
Equipment 40,586,000
Less: Accumulated depreciation-equipment 5,383,000 35,203,000
    Total plant assets £54,894,000


During 2017, the following selected cash transactions occurred.
Mar. 1 Purchased land for £2,052,000.
April 1 Sold equipment that cost £424,000 when purchased on January 1, 2013. The equipment was sold for £250,160.
June 1 Sold land purchased on June 1, 2007, for £1,533,000. The land cost £304,000.
Oct. 1 Purchased equipment for £1,498,000.
Dec. 31 Retired equipment that cost £503,000 when purchased on December 31, 2007. No residual value was received.
Journalize the above transactions. Bonita uses straight-line depreciation for buildings and equipment. The buildings are estimated to have a 50-year useful life and no residual value. The equipment is estimated to have a 10-year useful life and no residual value. Update depreciation on assets disposed of at the time of sale or retirement. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Mar. 1Apr. 1June 1Oct. 1Dec. 31

Mar. 1Apr. 1June 1Oct. 1Dec. 31

(To record depreciation)

Apr. 1

(To record sale of equipment)

Mar. 1Apr. 1June 1Oct. 1Dec. 31

Mar. 1Apr. 1June 1Oct. 1Dec. 31

Mar. 1Apr. 1June 1Oct. 1Dec. 31

(To record depreciation)

Mar. 1Apr. 1June 1Oct. 1Dec. 31

(To record retirement of equipment)
Answer #1
Date Account Titles and Explanation Debit Credit
Mar. 1 Land 2052000
Cash 2052000
(To record purchase of land)
Apr. 1 Depreciation expense 10600
Accumulated depreciation-equipment 10600
(To record depreciation)
Apr. 1 Cash 250160
Accumulated depreciation-equipment 180200
Gain on disposal 6360
Equipment 424000
(To record sale of equipment)
Jun. 1 Cash 1533000
Gain on disposal 1229000
Land 304000
(To record sale of land)
Oct. 1 Equipment 1498000
Cash 1498000
(To record purchase of equipment)
Dec. 31 Depreciation expense 50300
Accumulated depreciation-equipment 50300
(To record depreciation)
Dec. 31 Accumulated depreciation-equipment 503000
Equipment 503000
(To record retirement of equipment)

Working:

1. Annual depreciation on equipment = £424000/10 years = £42400

Depreciation for 3 months = £42400 x 3/12 = £10600

Accumulated depreciation = [(£42400 x 4 years) + £10600] = £169600 + £10600 = £180200

2. Annual depreciation on equipment = £503000/10 years = £50300